OpenStudy (anonymous):

Find the sum of a 10-term geometric sequence when the first term is 3 and the last term is 59,049 and select the correct answer below.

OpenStudy (anonymous):

177,147 88,572 88,575<---my answer 177,144

pooja195 (pooja195):

\[\huge~\rm~a_n=a_1\times~r^{n-1}\] \[\huge~\rm~a_n=3\times~r^{n-1} \] \[\huge~\rm~a_{10}=3\times~r^9 \] \[\huge~\rm~59049=3\times~r^9\] \[\huge~\rm~59049/3=r^9\]

pooja195 (pooja195):

\[\huge~\rm~19683=r^9\] \[\huge~\rm~\sqrt[9]{19683=r}\] From that we get r=3

OpenStudy (anonymous):

I got 2187? I think I did it wrong

pooja195 (pooja195):

Ye you are wrong :) look at the work above

OpenStudy (anonymous):

I dont know how to type that into my calculator :( I originally just did 19683 times 1/9

pooja195 (pooja195):

Well cacls wont be there all your life :(

OpenStudy (anonymous):

haha I suppose :)

pooja195 (pooja195):

\[\huge~\rm~\Large S_{10} = \frac{3 (1-3^{10})}{1-3} =88572\]

OpenStudy (anonymous):

:) So I was right?

pooja195 (pooja195):

You were off a few numbers look at your answer choices